OpenStudy (ketz):

Solve this quadratic equation: z2 + (2j - 3)z + (5 - j)=0

OpenStudy (anonymous):

is j here an imaginary number?

OpenStudy (anonymous):

if j is just imaginary number then just use quadratic formula where you coefficients are a=1, b=(2j-3) and c=(5-j) most likely the j will cancel out once you take root

OpenStudy (harsimran_hs4):

let z = x +jy substitute in the above eq and tell me what you get (do separate the real and complex terms when you report your answer)